Lilydale Ltd purchased new equipment on 1 January 2016, at a cost of $200,000. The company estimated that the equipment has a residual value of $20,000. The equipment is expected to be used for 5 years.

Required

Assuming the financial year ends on 31st Dec, calculate the depreciation expense using sum-of-years digit methods for years ended 2016 to 2019. Please show the calculation of the sum-of-years digit and depreciable amount. Note: The depreciation table is not required.
Answer #1

Sum of Years' digits = (1 + 2 + 3 + 4 + 5) = 15

Depreciable amount = (200,000 - 20,000) = 180,000

Depreciation expense = Depreciable amount * (Remaning useful life / Sum of Years' digits)

2016:
= Depreciation
= 180,000 * (5/15)
= 60,000

2017:
= Depreciation
= 180,000 * (4/15)
= 48,000

2018:
= Depreciation
= 180,000 * (3/15)
= 36,000

2019:
= Depreciation
= 180,000 * (2/15)
= 24,000

2020:
= Depreciation
= 180,000 * (1/15)
= 12,000